OpenStudy (anonymous):

For what x-value(s) does sin(x) = -1?

OpenStudy (anonymous):

\[\sin^{-1} (-1) = -\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

how would i find this?

OpenStudy (anonymous):

Either by using a calculator set to radian mode or memorize the unit circle. It's just a question of knowing your trig identities.

OpenStudy (anonymous):

ok sweet! thanks so much

OpenStudy (anonymous):

No prob, good luck!

OpenStudy (raden):

x = 270 degrees is correct too